OpenStudy (anonymous):

Simplify: t7 over t2

OpenStudy (solomonzelman):

\[Did~~~~you~~~~mean~~~~~t^7/t^2~~?\]

OpenStudy (solomonzelman):

when you divide you subtract exponents so \[t^7/t^2~=~t^{(7-2)}=t^5\]

undeadknight26 (undeadknight26):

what he said...

OpenStudy (solomonzelman):

Right, it's just the matter of knowing the rules of simplifying exponents.

OpenStudy (anonymous):

Yes Thank You you were correct!